A 3.47-? resistor and a 80.7-mF capacitor are connected in series and connected to a 11.9 V EMF. What is the time constant? How many time constants after the EMF is removed will it take to reach 1.23 V?

A 3.47-? resistor and a 80.7-mF capacitor are conn
Answer #1

Time Constant, t = R*C
t = 3.47*80.7*10^-3 s
t = 0.28 s
Time Constant, t = 0.28 s


q = Q*e^(-t/RC)
1.23 = 11.9 *e^(-t/0.28)
t = 0.635 s

No of time constant's = 0.635/0.28 = 2.26 Time constant

If they are asking in 1 sig figure, then 2 is the answer.